# Preference Robustness for DPO with Applications to Public Health

### v1

@@ Abstract
We study an LLM fine-tuning task for designing reward functions for sequential resource allocation problems in public health, guided by human preferences expressed in natural language. This setting presents a challenging testbed for alignment due to complex and ambiguous objectives and limited data availability. We propose DPO-PRO, a robust fine-tuning algorithm based on Direct Preference Optimization (DPO), which accounts for uncertainty in the preference distribution using a lightweight Distributionally Robust Optimization (DRO) formulation. Unlike prior DRO-based DPO methods, DPO-PRO is significantly less conservative. We evaluate DPO-PRO on a real-world maternal mobile health program operated by the non-profit organization ARMMAN, as well as on standard alignment benchmarks. Experimental results demonstrate that our method consistently improves robustness to noisy preference signals compared to existing DPO variants. Moreover, DPO-PRO achieves comparable performance to prior self-reflection-based baseline for reward function design, while requiring significantly lower inference-time cost.
@@ 1 Introduction
Limited resource allocation is a core challenge in public health, where decision-makers must prioritize individuals for care or intervention under tight financial and operational constraints. In maternal healthcare, for instance, preventative care awareness programs have been shown to significantly reduce maternal mortality (Helpmum 2025; Jahan et al. 2018; ARMMAN 2025), yet are often implemented by non-profit organizations that must serve large populations with limited resources. To support such efforts, it is crucial to develop allocation policies that make effective sequential decisions based on individual-level attributes such as age, income, or health status.
@@
Reinforcement learning (RL) offers a powerful framework for learning allocation policies in complex, dynamic environments. A widely used model in this context is the Restless Multi-Armed Bandit (RMAB) (Niño-Mora 2023), a class of Markov Decision Processes (MDPs) well-suited for sequential decision-making under resource constraints. However, in RL, it is well-recognized that designing an appropriate reward function is critical (Singh, Lewis, and Barto 2009; Booth et al. 2023). In public health applications, decision-makers’ priorities (e.g., which populations to prioritize) must be encoded into the reward function, so that resulting policy outcomes align with long-term public health goals. Crafting a reward function that accurately captures these priorities is often labor-intensive, demanding manual tuning, repeated simulation, and domain expertise. As public health goals shift over time, this design process must be repeated, limiting the scalability and adaptability of RL-based solutions in dynamic, real-world environments.
Figure 1: Overview of the implementation of DPO-PRO to public health applications such as the maternal health ARMMAN program in Section 5.2.
@@
Recently, several works explored using large language models (LLMs) to automate reward function design for sequential decision-making problems (Mirchandani, Karamcheti, and Sadigh 2021; Goyal, Niekum, and Mooney 2019; Carta et al. 2022; Ma et al. 2024; Behari et al. 2024; Verma et al. 2025). In these works, an LLM interprets a user’s high-level objective (e.g., desired policy outcomes) in natural language and outputs a reward function that guides RL policy training. This enables efficient workflow for developing human-aligned RL policies. Recent studies demonstrated that LLMs can effectively generate reward functions for RMAB problems in public health domains (Behari et al. 2024; Verma et al. 2025).
@@
Existing approaches often rely on inference-time optimization via self-reflection (Shinn et al. 2023), where the LLM iteratively refines its outputs based on simulated feedback. While this strategy offers flexibility, it incurs substantial inference-time costs due to repeated rounds of simulation and reasoning. Furthermore, in non-profit settings, data sensitivity constraints often limit the use of large proprietary models, leading organizations to adopt smaller open-source alternatives that may be less capable off-the-shelf but can be finetuned. Because self-reflection is highly dependent on model strength, its effectiveness may be limited in these contexts. Fine-tuning offers a more practical alternative by directly improving model capabilities while avoiding the high inference costs and computational overhead that become prohibitive in large-scale public health deployments (Wu et al. 2024), making it well-suited for resource-constrained domains such as public health (Zhao et al. 2025).
@@
Thus, we aim to move beyond inference-time methods and instead fine-tune an LLM to more effectively translate natural language preferences into high-performing reward functions for RMABs. A promising approach is Direct Preference Optimization (DPO) (Rafailov et al. 2023), which enables LLM fine-tuning from pairwise preference annotations. Recent work has also leveraged LLM-as-a-judge to scale annotation as a proxy for human feedback (Zhang et al. 2025; Tan et al. 2024; Cui et al. 2024; Zhu et al. 2024). However, applying these techniques to the public health domain introduces significant challenges, particularly due to the presence of noise in the preference signal.
@@
First, the objectives in the public health domain are often ambiguous or subjective (e.g., “prioritize older individuals”). Second, reward design in RMABs demands complex reasoning, as the quality of a proposed reward function is not evaluated in isolation, but rather by the policy it induces, increasing the complexity of assessing preference. Third, real-world public health datasets are typically smaller, increasing the risk of misalignment due to overfitting and reward hacking, which is a common problem with reinforcement learning for human feedback methods such as DPO (Gheshlaghi Azar et al. 2024; Amini, Vieira, and Cotterell 2024; Yang et al. 2024; Xiao et al. 2025). However, public health is a high-stakes domain, where poorly aligned policies can lead to serious real-world consequences.
Figure 2: DPO-PRO achieves comparable reward function performance to DLM on various ARMMAN maternal health tasks (Section 5.3) while providing (i) constant inference time scaling with only a finetuned LLama 8B LLM and (ii) removing external high-end LLM APIs for deployment.
@@
To tackle these challenges, we introduce DPO-PRO, a lightweight and distributionally robust enhancement of DPO (Rafailov et al. 2023). Its application to public health settings is summarized in Figure 1. DPO-PRO incorporates uncertainty in the preference distribution using an efficient yet effective Distributionally Robust Optimization (DRO) formulation (Rahimian and Mehrotra 2022), ensuring that the fine-tuned policy remains reliable under imperfect preference annotation. Unlike prior DRO-based DPO methods (Wu et al. 2025; Xu et al. 2025; Mandal, Sasnauskas, and Radanovic 2025), our approach avoids excessive conservatism and only adds negligible computational overhead. We further show that the resulting robust loss can be interpreted as a regularized DPO loss that penalizes model overconfidence and weak preference signals.
@@
We evaluate DPO-PRO in a simulated environment based on real-world maternal health application involving ARMMAN, a nonprofit organization that delivers preventative care messages via mobile health programs (ARMMAN 2025). Health workers can place a limited number of live service calls each week to boost beneficiary engagement. Prior work has modeled this problem using RMABs, where each beneficiary’s engagement evolves over time and the objective is to allocate $K$ live calls per week to maximize long-term engagement. Our experiments demonstrate that DPO-PRO is more robust to noisy preference supervision than both vanilla DPO and prior DRO-based approaches across a range of public health objectives. We also show that our method shows comparable performance with the Decision Language Model (DLM), a previous method based on self-reflection that was applied in the context of ARMMAN (Behari et al. 2024), but requires significantly lower inference-time cost as shown in Figure 2. Additionally, our finetuned LLM has the benefit of removing the need for external LLM APIs during deployment, which is required when working with sensitive or private data.
@@ 2 Related Works
A growing body of work has explored robustness in RLHF and DPO, particularly by leveraging DRO. Wu et al. (2025) introduce a DRO-based DPO formulation using a KL-divergence-based ambiguity set. Xu et al. (2025) extend this approach to both KL and Wasserstein distances, while Mandal, Sasnauskas, and Radanovic (2025) explore robustness under total variation distance, applying it to both DPO and standard two-stage RLHF.
In contrast to these DRO-based approaches, other works assume explicit corruption models. For example, Chowdhury, Kini, and Natarajan (2024) assume that preference labels are flipped according to a known noise rate, and Bukharin et al. (2024) model preferences using a Bradley–Terry (BT) framework (Bradley and Terry 1952) with added noise on the reward differences.
Hong et al. (2024) use a DRO framework to develop adaptive loss functions for reward model learning, but focus is not on distributional robustness. Zhan et al. (2024) introduce robustness by constructing confidence sets over learned reward functions, followed by pessimistic policy optimization.
Broadly, existing robustness techniques fall into two categories: (1) methods relying on strong assumptions about noise models, which may be unrealistic in practice, and (2) DRO-based methods that avoid explicit corruption modeling but hedge against worst-case distributions. As discussed in Section 3.2, standard DRO approaches tend to be overly conservative and often require solving challenging min–max optimization problems. For tractability, prior methods rely on heuristic approximations (Wu et al. 2025; Xu et al. 2025), potentially undermining theoretical guarantees or practical robustness. In contrast, we provide efficient approach DPO-PRO, with theoretical guarantees, and simultaneously illustrate its applicability in real world public health settings.
@@ 3.1 Data Distribution
Once the reward model is trained, a policy $\pi_{\theta}$ is optimized to maximize the expected reward while remaining close to a reference policy $\pi_{ref}$, typically via Proximal Policy Optimization (PPO) (Schulman et al. 2017):
@@
As discussed earlier, DPO is vulnerable to overfitting and noise in the data. To address this, recent work has proposed distributionally robust objectives of the form $\max_{P\in\mathcal{Q}(\mathcal{D})}\mathcal{L}_{{P}}(\pi_{\theta}),$ where $\mathcal{Q}(\mathcal{D})$ denotes an ambiguity set centered around the empirical distribution (data) $\mathcal{D}$ and the loss $\mathcal{L}_{{P}}$ is the DPO loss evaluated under the perturbed distribution $P$ rather than the original data distribution (Xu et al. 2025; Wu et al. 2025).
@@ 3.3 Reward Function Design Tasks for RMABs
In our setting, the prompt $x$ represents a desired policy outcome specified by a human prompter, for example, Focus on the young mothers by age and also focus on those with low income. See Figure 1. The responses $y_{1}$ and $y_{2}$ correspond to candidate reward functions for RMABs (details of the RMAB formulation are provided in the Appendix). The role of the LLM-annotator is to judge which of the two reward functions, $y_{1}$ or $y_{2}$, better aligns with the objective expressed in the prompt.
A key challenge in this task is that the quality of a reward function cannot be assessed in isolation. Rather, it must be evaluated based on the policy it induces when optimized, i.e., how effectively the resulting policy fulfills the prompter’s (often ambiguous) intent. This requires the annotator to infer the downstream effects of each reward function. Given the complexity and ambiguity involved, the resulting preference signals are inevitably noisy and should not be treated as fully reliable.
@@ 4 DPO with Preference Robustness
In this section, we introduce DPO-PRO, a distributionally robust version of DPO that specifically targets distributional shifts or other forms of noise in the preference distribution.
@@ 4.1 Uncertainties in the Preference Distribution
In each data point $(x,y_{1},y_{2},c)$, the triplet $(x,y_{1},y_{2})$ represents observed content, which typically comes from a well-understood and controllable data collection process. In general, $x$ is drawn from a pool of candidate inputs, and the response pair $(y_{1},y_{2})$ is generated by the reference policy $\pi_{ref}$ (usually a supervised-fine tuned model) for every sampled prompt $x$. As more data is collected, the uncertainty in this part of the data is expected to diminish. Noise in the prompts and responses becomes even less of an issue under iterative DPO methods, which are increasingly adopted in practice (Xiong et al. 2024; Cen et al. 2025; Xie et al. 2025).
@@
where $Q(x,y_{1},y_{2}\rho)=\left\{p:\chi^{2}\big{(}p\,\|\,q(y_{1}\succ y_{2}\mid x)\big{)}\leq\rho\right\}$. Our formulation avoids unnecessary robustification over the joint distribution on $(x,y_{1},y_{2})$, and assumes that this part of the data is reliable. Furthermore, unlike standard DRO methods that can add significant computational burden (Rahimian and Mehrotra 2022), our formulation introduces negligible additional cost as shown next.
@@ 4.3 Efficient Optimization of the Loss
Note that we do not formally differentiate through the inner maximization in Eq (1) with respect to $\theta$. However, Danskin’s theorem (Bertsekas 1999) justifies our approach. For a fixed $\theta$, we may solve the inner maximization and directly substitute the resulting worst-case $\hat{p}$ into the gradient expression.
@@ 4.5 Obtaining the Soft Score $q$ in Practice
The DRO adjustment described above requires a soft score $q(y_{1}\succ y_{2}|x)$, rather than a binary annotation for each prompt and response pair $(x,y_{1},y_{2})$. In this work we focus on settings where preferences are evaluated by LLM judges. Soft scores can be obtained from LLM judges in several ways: (1) producing multiple binary preference judgments through repeated queries and averaging the outcomes; (2) directly outputting a numerical preference score via appropriate prompting; or (3) extracting the log-probabilities assigned to tokens representing the preferences (e.g., 1 for $y_{1}\succ y_{2}$ and -1 for $y_{2}\succ y_{1}$), and applying a softmax transformation to estimate a preference distribution (Lee et al. 2024). We provide further discussion on annotation formats and the connection to robustness in the Appendix.
@@ 5 Experiments
We validate DPO-PRO in two different settings. First, we validate the methodology on a standardized benchmark for preference-based LLM finetuning. Next, we adapt this experiment for the ARMMAN maternal health dataset (ARMMAN 2025) in the following section.

### v2

@@ Abstract
We study an LLM fine-tuning task for designing reward functions for sequential resource allocation problems in public health, guided by human preferences expressed in natural language. This setting presents a challenging testbed for alignment due to complex and ambiguous objectives and limited data availability. We propose DPO-PRO, a robust fine-tuning algorithm based on Direct Preference Optimization (DPO), which accounts for uncertainty in the preference distribution using a lightweight Distributionally Robust Optimization (DRO) formulation. Unlike prior DRO-based variants, DPO-PRO focuses solely on uncertainty in preferences, avoiding unnecessary conservatism and incurring negligible computational overhead. We evaluate DPO-PRO on a real-world maternal mobile health program operated by the non-profit organization ARMMAN, as well as on standard alignment benchmarks. Experimental results demonstrate that our method consistently improves robustness to noisy preference signals compared to existing DPO variants. Moreover, DPO-PRO achieves comparable performance to prior self-reflection-based baseline for reward function design, while requiring significantly lower inference-time cost.
@@ 1 Introduction
Limited resource allocation is a core challenge in public health, where decision-makers must prioritize individuals for care or intervention under tight financial and operational constraints. In maternal healthcare, for instance, preventative care awareness programs have been shown to significantly reduce maternal mortality (Helpmum; Jahan2018MomConnect; ARMMAN), yet are often implemented by non-profit organizations that must serve large populations with limited resources. To support such efforts, it is crucial to develop allocation policies that make effective sequential decisions based on individual-level attributes such as age, income, or health status.
@@
Reinforcement learning (RL) offers a powerful framework for learning allocation policies in complex, dynamic environments. A widely used model in this context is the Restless Multi-Armed Bandit (RMAB) (math11071639), a class of Markov Decision Processes (MDPs) well-suited for sequential decision-making under resource constraints. However, in RL, it is well-recognized that designing an appropriate reward function is critical (singh09; Booth_Knox_Shah_Niekum_Stone_Allievi_2023). In public health applications, decision-makers’ priorities (e.g., which populations to prioritize) must be encoded into the reward function, so that resulting policy outcomes align with long-term public health goals. Crafting a reward function that accurately captures these priorities is often labor-intensive, demanding manual tuning, repeated simulation, and domain expertise. As public health goals shift over time, this design process must be repeated, limiting the scalability and adaptability of RL-based solutions in dynamic, real-world environments.
@@
Recently, several works explored using large language models (LLMs) to automate reward function design for sequential decision-making problems (mirchan21; ijcai2019p331; carta2022eager; ma2024eureka; behari2024a; verma2025balancingactprioritizationstrategies). In these works, an LLM interprets a user’s high-level objective (e.g., desired policy outcomes) in natural language and outputs a reward function that guides RL policy training. This enables efficient workflow for developing human-aligned RL policies. Recent studies demonstrated that LLMs can effectively generate reward functions for RMAB problems in public health domains (behari2024a; verma2025balancingactprioritizationstrategies).
@@
Existing approaches often rely on inference-time optimization via self-reflection (shinn2023reflexion), where the LLM iteratively refines its outputs based on simulated feedback. Although this strategy provides flexibility, it incurs substantial inference-time costs due to repeated rounds of simulation and reasoning. In many non-profit settings, data sensitivity constraints further limit the use of large proprietary models, pushing organizations toward open-source alternatives, which are often smaller and less capable. Because self-reflection is highly dependent on model strength, its effectiveness can be diminished in these environments. As a result, fine-tuning open-source models offers a more practical path: it directly improves their capabilities while avoiding the high inference costs and privacy issues in large-scale public health deployments (wu2024inference). This makes fine-tuning particularly suitable for resource-constrained domains such as public health (zhao2025foundation).
@@
Thus, we aim to move beyond inference-time methods and instead fine-tune an LLM to more effectively translate natural language preferences into high-performing reward functions for RMABs. A promising approach is Direct Preference Optimization (DPO) (rafailov2023DPO), which enables LLM fine-tuning from pairwise preference annotations. Recent work has also leveraged LLM-as-a-judge to scale annotation as a proxy for human feedback (zhang2025aaai; tan-etal-2024-large; ultra2024; zhu2024starlingb). However, applying these techniques to the public health domain introduces significant challenges, particularly due to the presence of noise in the preference signal.
@@
First, the objectives in the public health domain are often ambiguous or subjective (e.g., “prioritize older individuals”). Second, reward design in RMABs demands complex reasoning, as the quality of a proposed reward function is not evaluated in isolation, but rather by the policy it induces, increasing the complexity of assessing preference. Third, real-world public health datasets are typically smaller, increasing the risk of misalignment due to overfitting and reward hacking, which is a common problem with reinforcement learning for human feedback methods such as DPO (pmlr-v238-gheshlaghi-azar24a; amini-etal-2024-direct; yang2024orthogonalfinetuningdirectpreference; xiao2025comprehensivesurveydirectpreference). However, public health is a high-stakes domain, where poorly aligned policies can lead to serious real-world consequences.
@@
To tackle these challenges, we introduce DPO-PRO, a lightweight and distributionally robust enhancement of DPO. DPO-PRO incorporates uncertainty in the preference distribution using an efficient yet effective Distributionally Robust Optimization (DRO) formulation (Rahimian_2022), ensuring that the fine-tuned policy remains reliable under imperfect preference annotation. Unlike prior DRO-based DPO methods (wu2025drdpo; xu2025distributionallyrobustdirectpreference; mandal2025distributionallyrobustreinforcementlearning), our approach avoids excessive conservatism and only adds negligible computational overhead. We further show that the resulting robust loss can be interpreted as a regularized DPO loss that penalizes model overconfidence and weak preference signals.
@@
We evaluate DPO-PRO on the UltraFeedback benchmark (ultra2024) and on a simulated maternal health task based on ARMMAN’s real-world mobile health program (ARMMAN). In this setting, health workers can place only a limited number of live service calls each week to boost beneficiary engagement. Prior work models this problem using RMABs, where each beneficiary’s engagement evolves over time and the objective is to allocate a limited number of calls to maximize long-term engagement. Our experiments show that DPO-PRO is more robust to noisy preference annotation than vanilla DPO and prior DRO-based approaches, both on UltraFeedback and across a range of public health objectives. We further show that DPO-PRO achieves performance comparable to the Decision Language Model (DLM), a self-reflection–based method used in the ARMMAN setting (behari2024a), while requiring significantly lower inference-time cost (See Figure 1 in Section 5.2).
@@ 2 Related Works
A growing body of work has explored robustness in RLHF and DPO, particularly by leveraging DRO. wu2025drdpo introduce a DRO-based DPO formulation using a KL-divergence-based ambiguity set. xu2025distributionallyrobustdirectpreference extend this approach to both KL and Wasserstein distances, while mandal2025distributionallyrobustreinforcementlearning explore robustness under total variation distance, applying it to both DPO and standard two-stage RLHF.
In contrast to these DRO-based approaches, other works assume explicit corruption models. For example, chowdhury2024 assume that preference labels are flipped according to a known noise rate, and bukharin2024robust model preferences using a Bradley–Terry (BT) framework (BT) with added noise on the reward differences.
hong2024adaptive use a DRO framework to develop adaptive loss functions for reward model learning, but focus is not on distributional robustness. zhan2024provable introduce robustness by constructing confidence sets over learned reward functions, followed by pessimistic policy optimization.
Broadly, existing robustness techniques fall into two categories: (1) methods relying on strong assumptions about noise models, which may be unrealistic in practice, and (2) DRO-based methods that avoid explicit corruption modeling but hedge against worst-case distributions. As discussed in Section 3.2, standard DRO approaches tend to be overly conservative and often require solving challenging min–max optimization problems. For tractability, prior methods rely on heuristic approximations (wu2025drdpo; xu2025distributionallyrobustdirectpreference), potentially undermining theoretical guarantees or practical robustness. In contrast, we provide efficient approach DPO-PRO, with theoretical guarantees, and simultaneously illustrate its applicability in real world public health settings.
@@ 3.1 Data Distribution
Once the reward model is trained, a policy $\pi_{\theta}$ is optimized to maximize the expected reward while remaining close to a reference policy $\pi_{ref}$, typically via Proximal Policy Optimization (PPO) (schulman2017ppo):
@@
As discussed earlier, DPO is vulnerable to overfitting and noise in the data. To address this, recent work has proposed distributionally robust objectives of the form $\max_{P\in\mathcal{Q}(\mathcal{D})}\mathcal{L}_{{P}}(\pi_{\theta}),$ where $\mathcal{Q}(\mathcal{D})$ denotes an ambiguity set centered around the empirical distribution (data) $\mathcal{D}$ and the loss $\mathcal{L}_{{P}}$ is the DPO loss evaluated under the perturbed distribution $P$ rather than the original data distribution (xu2025distributionallyrobustdirectpreference; wu2025drdpo).
@@ 3.3 Reward Function Design Tasks for RMABs
In our setting, the prompt $x$ represents a desired policy outcome specified by a human prompter, for example, Focus on the young mothers by age and also focus on those with low income. The responses $y_{1}$ and $y_{2}$ correspond to candidate reward functions for RMABs. The role of the LLM-annotator is to judge which of the two reward functions, $y_{1}$ or $y_{2}$, better aligns with the objective expressed in the prompt.
A key challenge in this task is that the quality of a reward function cannot be assessed in isolation. Rather, it must be evaluated based on the policy it induces when optimized, i.e., how effectively the resulting policy fulfills the user’s (often ambiguous) intent. This requires the annotator to infer the downstream effects of each reward function. Given the complexity and ambiguity involved, the resulting preference signals are inevitably noisy and should not be treated as fully reliable.
@@ 4 DPO with Preference Robustness
In this section, we introduce DPO-PRO, a distributionally robust version of DPO that specifically targets distributional shifts or other forms of noise in the preference distribution. The complete proofs for all theorems are provided in the extended version of this paper (kim2025dpopro), which includes the full technical appendix.
@@ 4.1 Uncertainties in the Preference Distribution
In each data point $(x,y_{1},y_{2},c)$, the triplet $(x,y_{1},y_{2})$ represents observed content, which typically comes from a well-understood and controllable data collection process. In general, $x$ is drawn from a pool of candidate inputs, and the response pair $(y_{1},y_{2})$ is generated by the reference policy $\pi_{ref}$ (usually a supervised-fine tuned model) for every sampled prompt $x$. As more data is collected, the uncertainty in this part of the data is expected to diminish. Noise in the prompts and responses becomes even less of an issue under iterative DPO methods, which are increasingly adopted in practice (xiong2024iterative; cen2025valueincentivized; xie2025exploratory).
@@
where $Q(x,y_{1},y_{2}\rho)=\left\{p:\chi^{2}\big(p\,\|\,q(y_{1}\succ y_{2}\mid x)\big)\leq\rho\right\}$. Our formulation avoids unnecessary robustification over the joint distribution on $(x,y_{1},y_{2})$, and assumes that this part of the data is reliable. Furthermore, unlike standard DRO methods that can add significant computational burden (Rahimian_2022), our formulation introduces negligible additional cost as shown next.
@@ 4.3 Efficient Optimization of the Loss
Note that we do not formally differentiate through the inner maximization in Eq (1) with respect to $\theta$. However, Danskin’s theorem (Bertsekas99Np) justifies our approach. For a fixed $\theta$, we may solve the inner maximization and directly substitute the resulting worst-case $\hat{p}$ into the gradient expression.
@@ 4.5 Obtaining the Soft Score $q$ in Practice
The DRO adjustment described above requires a soft score $q(y_{1}\succ y_{2}|x)$, rather than a binary annotation for each prompt and response pair $(x,y_{1},y_{2})$. Soft scores can be obtained from LLM judges in several ways: (1) producing multiple binary preference judgments through repeated queries and averaging the outcomes; (2) directly outputting a numerical preference score via appropriate prompting; or (3) extracting the log-probabilities assigned to tokens representing the preferences (e.g., 1 for $y_{1}\succ y_{2}$ and -1 for $y_{2}\succ y_{1}$), and applying a softmax transformation to estimate a preference distribution (lee2024). When using a off-the-shelf reward model, the scalar outputs of the reward model can be converted into a pairwise preference probability using the BT model. We provide further discussion on annotation formats and the connection to robustness in (kim2025dpopro).
@@ 5 Experiments
We validate DPO-PRO in two different settings. First, we validate the methodology on a standardized benchmark for preference-based LLM finetuning. Next, we adapt this experiment for the ARMMAN maternal health dataset (ARMMAN) in the following section. The complete experimental details are provided in the extended version of this paper (kim2025dpopro).